Apr 1, 2018

From Hall A Wiki
Jump to: navigation, search

Back to Main Page >> Moller Page >> Moller Weekly Meeting Index

previous meeting << >> following meeting


Moller DocDB

Original Moller Polarimeter Page

Moller ELog


Agenda:

  • Commissioning status
  • Simulation update -- Eric
  • Simulation update -- Bill

Connection details:

Meeting URL https://bluejeans.com/388466836

Meeting ID 388 466 836

US toll free phone 1.888.240.2560

Dial in, type in the meeting ID and then press ##


Minutes


Simona reported on data from commissioning when the foil angle was changed. Something is strange in the data that shows the angle is not reliable. The requested angle which relies on counting motor steps does not give the same angle as the encoder readback. This issue needs to be resolved. The plan is to remove the motion mechanism and directly test it outside the beam pipe. This should happen later this week. The target ladder will be given to Dave Meekins for storage. Simona has offered to give a few slides on what is included in the datastream for the Moller DAQ. We have requested that Dave Meekins install a high purity 10 micron copper foil in the remaining empty target position. We should hear back soon on whether or not this is possible in the short time we have at our disposal.

Eric showed a preliminary analysis trying to reconcile rate plots for the Quad scans. To be continued, but many questions.